# Qi Sun1,2*, Edoardo Cetin1*, Yujin Tang1*

### v2

@@ Abstract
Self-adaptive large language models (LLMs) aim to solve the challenges posed by traditional fine-tuning methods, which are often computationally intensive and static in their ability to handle diverse tasks. We introduce $\text{Transformer}^{2}$, a novel self-adaptation framework that adapts LLMs for unseen tasks in real-time by selectively adjusting only the singular components of their weight matrices. During inference, $\text{Transformer}^{2}$ employs a two-pass mechanism: first, a dispatch system identifies the task properties, and then task-specific “expert” vectors, trained using reinforcement learning, are dynamically mixed to obtain targeted behavior for the incoming prompt. Our method outperforms ubiquitous approaches such as LoRA, with fewer parameters and greater efficiency. $\text{Transformer}^{2}$ demonstrates versatility across different LLM architectures and modalities, including vision-language tasks. $\text{Transformer}^{2}$ represents a significant leap forward, offering a scalable, efficient solution for enhancing the adaptability and task-specific performance of LLMs, paving the way for truly dynamic, self-organizing AI systems. Our code is available at https://github.com/SakanaAI/self-adaptive-llms
@@ 1 Introduction
Self-adaptive large language models (LLMs) would represent a significant advancement in artificial intelligence, providing a framework where models can adjust to varied tasks and dynamic contexts in real time. While compositionality and scalability are crucial for effective adaptation, current LLM training methodologies fall short of achieving both these properties simultaneously. Our research aims to present a pioneering solution to realize this vision and address these gaps.
@@
We then introduce our full $\text{Transformer}^{2}$ framework to empower LLMs through the underlying principles of self-adaptation. Given a prompt from an unknown task, $\text{Transformer}^{2}$ entails a two-pass inference mechanism which we illustrate in Figure 1. During the first pass, $\text{Transformer}^{2}$ executes the model and observes its test-time behavior, gathering the relevant information to understand the necessary skills to tackle the current problem. During the second pass, our framework uses this information to combine the available expert vectors and provide a new modification to the base weights of the LLM specifically tailored to its test-time conditions. We design three different adaptation strategies that can be used within $\text{Transformer}^{2}$, which we show provide monotonic performance benefits with increasing access to the test-time conditions.
@@ 2 Related works
Self-adaptive LLMs We define self-adaptive LLMs as a group of LLMs or a standalone LLM that can evaluate and modify its behavior in response to changes in its operating environment or internal state, without external intervention. This adaptation can be explored from two perspectives: a macroview, where multiple LLMs collaborate and/or compete, and a microview, where internal adaptations allow a single LLM to specialize in different tasks.
@@
Low-rank adaptation PEFT methods such as LoRA (Hu et al., 2021) works by freezing the original model’s parameters and introducing small trainable low-rank matrices for task-specific updates. It significantly lowers the computational and memory costs while providing performance comparable to full fine-tuning. Inspired by LoRA’s design, various modifications have been proposed (Zhang et al., 2023; Kopiczko et al., 2023; Liu et al., 2024; Bałazy et al., 2024; Cetoli, 2024). $\text{Transformer}^{2}$ does not rely on low-rank matrices, and instead scales the singular vectors of the original parameter matrix that span the full rank space.
SVD for LLM Fine-tuning SVD is increasingly being used as an inductive bias for PEFT in LLMs. For example, Wang et al. (2024) decompose a weight matrix and use the minor singular components, associated with noisy or long-tail information, to initialize low-rank matrices for LoRA fine-tuning. In a similar vein, SVD is employed to approximate an original weight matrix with the top $r$ singular vectors, corresponding to the highest singular values. A small trainable matrix is then introduced on top of the truncated singular value matrix to adjust the magnitude and orientations within this top-$r$ subspace (Bałazy et al., 2024; Cetoli, 2024). However, the drawback of this approach is that retaining only the top singular components can result in the loss of important information, particularly when the singular values distribution is less skewed. The work most similar to ours is a concurrent effort by Lingam et al. (2024), where they introduce various sparsification methods that utilize the SVD of the weights. However, it is not for self-adaptive LLMs and does not use RL to enhance learning efficiency.
@@ 5 Conclusion
6 Author contributions
Yujin Tang initiated the project. Qi Sun proposed the prompted-based method, developed the evaluation framework, conducted the experiments, and provided contributions to writing. Edoardo Cetin designed the few-shot CEM adaptation strategy, performed the experiment, and made major contributions to manuscript writing. Yujin Tang proposed the core algorithm, conducted initial experiments, made major contributions to the manuscript, and managed the project.

### v3

# Transformer-Squared: Self-adaptive LLMs

@@ Abstract
Self-adaptive large language models (LLMs) aim to solve the challenges posed by traditional fine-tuning methods, which are often computationally intensive and static in their ability to handle diverse tasks. We introduce $\text{Transformer}^{2}$ (Transformer-Squared), a novel self-adaptation framework that adapts LLMs for unseen tasks in real-time by selectively adjusting only the singular components of their weight matrices. During inference, $\text{Transformer}^{2}$ employs a two-pass mechanism: first, a dispatch system identifies the task properties, and then task-specific “expert” vectors, trained using reinforcement learning, are dynamically mixed to obtain targeted behavior for the incoming prompt. Our method consistently outperforms ubiquitous approaches such as LoRA, with fewer parameters and greater efficiency. Furthermore, $\text{Transformer}^{2}$ demonstrates versatility across different LLM architectures and modalities, including vision-language tasks. $\text{Transformer}^{2}$ represents a significant leap forward, offering a scalable, efficient solution for enhancing the adaptability and task-specific performance of LLMs, paving the way for truly dynamic, self-organizing AI systems. We provide our full source code at https://github.com/SakanaAI/self-adaptive-llms.
@@ 1 Introduction
Self-adaptive large language models (LLMs) would represent a significant advancement in artificial intelligence, providing a framework where models can adjust to varied tasks and dynamic contexts in real time. This concept draws inspiration from the long-standing idea of neural networks modifying their own weights to adapt to tasks dynamically (Schmidhuber, 1993; Irie et al., 2022) and neural networks generating weights for other networks, as popularized by HyperNetworks and related methods (Ha et al., 2017; Stanley et al., 2009). While compositionality and scalability are crucial for effective adaptation, current LLM training methodologies fall short of achieving both these properties simultaneously. Our research aims to present a pioneering solution to realize this vision.
@@
We then introduce our full $\text{Transformer}^{2}$ (Transformer-Squared) framework to empower LLMs through the underlying principles of self-adaptation. Given a prompt from an unknown task, $\text{Transformer}^{2}$ entails a two-pass inference mechanism which we illustrate in Figure 1. During the first pass, $\text{Transformer}^{2}$ executes the model and observes its test-time behavior, gathering the relevant information to understand the necessary skills to tackle the current problem. During the second pass, our framework uses this information to combine the available expert vectors and provide a new modification to the base weights of the LLM specifically tailored to its test-time conditions. We design three different adaptation strategies that can be used within $\text{Transformer}^{2}$, which we show provide monotonic performance benefits with increasing access to the test-time conditions.
@@ 2 Related works
Self-adaptive LLMs We define self-adaptive LLMs as a group of LLMs or a standalone LLM that can evaluate and modify its behavior in response to changes in its operating environment or internal state, without external intervention. This dynamic adjustment has parallels to concepts like fast-weight memories, which enable networks to update weights in response to task demands (Schmidhuber, 1992; Gomez & Schmidhuber, 2005), and neural network weights being treated as dynamic programs (Schmidhuber, 2015). Recently, Panigrahi et al. (2023) introduces an approach where a smaller auxiliary transformer is updated dynamically within a larger model, aligning with the principles of self-adaptive behavior.
This adaptation can be explored from two perspectives: a macroview, where multiple LLMs collaborate and/or compete, and a microview, where internal adaptations allow a single LLM to specialize in different tasks.
@@
Low-rank adaptation PEFT methods such as LoRA (Hu et al., 2021) works by freezing the original model’s parameters and introducing small trainable low-rank matrices for task-specific updates. It significantly lowers the computational and memory costs while providing performance comparable to full fine-tuning. Inspired by LoRA’s design, various modifications have been proposed (Zhang et al., 2023; Kopiczko et al., 2023; Liu et al., 2024; Bałazy et al., 2024; Cetoli, 2024; anonymous2025eigenlora). $\text{Transformer}^{2}$ does not rely on low-rank matrices, and instead scales the singular vectors of the original parameter matrix that span the full rank space.
SVD for LLM Fine-tuning SVD is increasingly being used as an inductive bias for PEFT in LLMs. For example, Wang et al. (2024) decompose a weight matrix and use the minor singular components, associated with noisy or long-tail information, to initialize low-rank matrices for LoRA fine-tuning. Earlier work proposed using compressed forms like DCT coefficients for generating weight matrices in neural networks (Koutnik et al., 2010), offering efficiency in memory-constrained environments, which resonates with our approach. In a similar vein, SVD is employed to approximate an original weight matrix with the top $r$ singular vectors, corresponding to the highest singular values. A small trainable matrix is then introduced on top of the truncated singular value matrix to adjust the magnitude and orientations within this top-$r$ subspace (Bałazy et al., 2024; Cetoli, 2024). However, the drawback of this approach is that retaining only the top singular components can result in the loss of important information, particularly when the singular values distribution is less skewed. The work most similar to ours is a concurrent effort by Lingam et al. (2024), where they introduce various sparsification methods that utilize the SVD of the weights. However, it is not for self-adaptive LLMs and does not use RL to enhance learning efficiency.